# arXiv:2512.12436v2[cs.LG] 27 Feb 2026

### v2

@@ 3.3 Theoretical Background
$S=\begin{pmatrix}S_{[1:N_{1}]}&\mathbf{0}&\mathbf{0}&\dots&\mathbf{0}\\ \mathbf{0}&S_{[N_{1}+1:N_{2}]}&\mathbf{0}&\dots&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&S_{[N_{2}+1:N_{3}]}&\dots&\mathbf{0}\\ \dots&\dots&\dots&\dots&\dots\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\dots&S_{[N_{k-1}+1:N_{k}]}\end{pmatrix}$, (23)
whereby the symmetric matrices $S_{[N_{i-1}+1:N_{i}]}$, $i=1,\dots,k$ are the similarity matrices within the clusters $C_{1},\dots,C_{k}$ (with zero diagonals), $N_{0}=0.N_{i}=N_{i-1}+n_{i}$ for $i\geq 1$, $0<n_{1}<n_{2}<\dots<n_{k}$, $n_{k}$ being the number of objects.
Let $D(S)$ be the diagonal matrix such that $d_{j,j}=\sum_{\ell=1}^{n_{k}}s_{j,\ell}$ and $L(S)$ be the Laplacian L(S)=D(S)SL(S)=D(S)-S.
@@
$S^{\prime}=\begin{pmatrix}S_{[1:N_{1}]}&\mathbf{0}&\mathbf{0}&\dots&\mathbf{0}&0\\ \mathbf{0}&S_{[N_{1}+1:N_{2}]}&\mathbf{0}&\dots&\mathbf{0}&0\\ \mathbf{0}&\mathbf{0}&S_{[N_{2}+1:N_{3}]}&\dots&\mathbf{0}&0\\ \dots&\dots&\dots&\dots&\dots&\dots\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\dots&S_{[N_{k-1}+1:N_{k}]}&0\\ 0&0&0&\dots&0&0\end{pmatrix}$, (24)
@@
This may happen as follows: In case we take the eigenvector $\mathbf{v}_{k+1}$ for embedding the objects, then one of the intrinsic clusters will have the center at $\mathbf{0}$ (coordinate origin), and the other at $(0,...,1/\sqrt{n_{i}},....0)^{T}$. So the distance of the dissimilar object to the intrinsic cluster centers will amount either to $1$ or to $\sqrt{1+\frac{1}{n_{i}}}$ whereas the distance between intrinsic clusters will amount to either $1/\sqrt{n_{i}}$ or to $\sqrt{\frac{1}{n_{i^{\prime}}}+\frac{1}{n_{i}}}$. With increasing cluster sizes the distance between true clusters will decrease down to zero, while the dissimilar element’s distance will get closer to 1. At a point it will be more advantageous for $k$-means to separate the dissimilar element as a single cluster and to collapse two intrinsic clusters. With increasing number of elements dissimilar to all the other elements, the situation will get worse.
@@ 4.3 Real data
This fact causes the $L$-based and $K$-based and $\mathcal{B}$-based clustering algorithms to perform poorly, see Table 4, 3, 6. Compared to them, the $N$-based algorithm performs quite well, see Table 5.
@@
The tables 7, 8, 10, 9 show the impact of this filtering. In each case one can see an improvement of the performance.
@@ 4.5 Real Data Experiments - Multiple Clustering Methods
The Tables 13, …, 18 present results of our experiments about the impact on explainability when performing rough-set based removal of boundary data.

### v3

# Rough Sets for Explainability of Spectral Graph Clustering

@@ 3.3 Theoretical Background
$S=\begin{pmatrix}S_{[[1]]}&\mathbf{0}&\mathbf{0}&\dots&\mathbf{0}\\ \mathbf{0}&S_{[[2]]}&\mathbf{0}&\dots&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&S_{[[3]]}&\dots&\mathbf{0}\\ \dots&\dots&\dots&\dots&\dots\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\dots&S_{[[k]]}\end{pmatrix}$, (23)
whereby the symmetric matrices $S_{[[i]]}$, $i=1,\dots,k$ are the similarity matrices within the clusters $C_{1},\dots,C_{k}$ (with zero diagonals), with the number of objects within the cluster $C_{i}$ equal to $n_{i}>0$. Let further $N_{0}=0$ and $N_{i}=N_{i-1}+n_{i}$ for $i=1,\dots,k$.
Let $D(S)$ be the diagonal matrix such that $d_{j,j}=\sum_{\ell=1}^{n_{k}}s_{j,\ell}$ and $L(S)$ be the combinatorial Laplacian L(S)=D(S)SL(S)=D(S)-S.
$L(S)=\begin{pmatrix}L(S_{[[1]]})&\mathbf{0}&\mathbf{0}&\dots&\mathbf{0}\\ \mathbf{0}&L(S_{[[2]]})&\mathbf{0}&\dots&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&L(S_{[[3]]})&\dots&\mathbf{0}\\ \dots&\dots&\dots&\dots&\dots\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\dots&L(S_{[[k]]})\end{pmatrix}$, (24)
@@
$S^{*}=\begin{pmatrix}S&\mathbf{0}\\ \mathbf{0}&0\end{pmatrix}=\begin{pmatrix}S_{[[1]]}&\mathbf{0}&\mathbf{0}&\dots&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&S_{[[2]]}&\mathbf{0}&\dots&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&S_{[[3]]}&\dots&\mathbf{0}&\mathbf{0}\\ \dots&\dots&\dots&\dots&\dots&\dots\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\dots&S_{[[k]]}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\dots&\mathbf{0}&0\end{pmatrix}$, (25)
Then
$L(S^{*})=L\begin{pmatrix}S&\mathbf{0}\\ \mathbf{0}&0\end{pmatrix}=\begin{pmatrix}L(S_{[[1]]})&\mathbf{0}&\mathbf{0}&\dots&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&L(S_{[[2]]})&\mathbf{0}&\dots&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&L(S_{[[3]]})&\dots&\mathbf{0}&\mathbf{0}\\ \dots&\dots&\dots&\dots&\dots&\dots\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\dots&L(S_{[[k]]})&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\dots&\mathbf{0}&0\end{pmatrix}$, (26)
@@
This may happen as follows: Consider two datapoints $j$ and $j^{\prime}$ belonging to distinct intrinsic clusters $C_{i}$ and $C_{i^{\prime}}$. The vector difference between them is equal $o_{j}-o_{j^{\prime}}=(0,\dots,1/\sqrt{n_{i}},0,\dots,-1/\sqrt{n_{i^{\prime}}},0,\dots,0)^{T}$. Then length of $o_{j}-o_{j^{\prime}}$ will be equal to $\sqrt{\frac{1}{n_{i^{\prime}}}+\frac{1}{n_{i}}}$. On the other hand, the vector difference $o_{j}-o_{N_{k}+1}$ them is equal $o_{j}-o_{N_{k}+1}=(0,\dots,1/\sqrt{n_{i}},0,\dots,-1,0,\dots,0)^{T}$. Then length of $o_{j}-o_{N_{k}+1}$ will be equal to $\sqrt{1+\frac{1}{n_{i}}}$. The table 1 presents example eigenvectors obtained when adding 1 or 4 unrelated documents to a document collection with cleanly separated 4 clusters of 2,3,4,5 documents resp.
With increasing cluster sizes the distance between true clusters $\|o_{j}-o_{j^{\prime}}\|$ will decrease down to zero, while the dissimilar element’s distance to objects of any cluster $\|o_{j}-o_{N_{k}+1}\|$ will get closer to 1. At a point it will be more advantageous for $k$-means to separate the dissimilar element as a single cluster and to collapse two intrinsic clusters. With increasing number of elements dissimilar to all the other elements, the situation will get worse.
@@
To check this, we generated a similarity matrix for 4 clusters with cluster sizes 8 12 16 20. Then added noise in the range from=0., to=0.000001. $k$-means (with tenfold restart) applied to the space spanned by 4 eigenvectors of the combinatorial Laplacian related to 4 lowest eigenvalues retrieved correctly 4 clusters with sizes 8 12 16 20.
However, for the same design, if one unrelated document was added, one got cluster sizes 1, 16, 28, 12. With two unrelated documents, cluster sizes were 8 12 16 20 . With 4 unrelated documents the cluster sizes were 17, 21, 21, 1.
For true clusters of sizes 16 24 32 40 the returned clusters, when 4 unrelated documents were added, were: 97, 2, 1, 16
$\mathbf{o}$ $\mathbf{v}_{1}$ $\mathbf{v}_{2}$ $\mathbf{v}_{3}$ $\mathbf{v}_{5}$ $\mathbf{v}_{4}$
1 0 0 0.707 0 0
2 0 0 0.707 0 0
3 0 0 0 0 -0.577
4 0 0 0 0 -0.577
5 0 0 0 0 -0.577
6 0 -0.5 0 0 0
7 0 -0.5 0 0 0
8 0 -0.5 0 0 0
9 0 -0.5 0 0 0
10 -0.447 0 0 0 0
11 -0.447 0 0 0 0
12 -0.447 0 0 0 0
13 -0.447 0 0 0 0
14 -0.447 0 0 0 0
15 0 0 0 1 0
$\mathbf{o}$ $\mathbf{v}_{1}$ $\mathbf{v}_{2}$ $\mathbf{v}_{3}$ $\mathbf{v}_{5}$ $\mathbf{v}_{6}$ $\mathbf{v}_{7}$ $\mathbf{v}_{8}$ $\mathbf{v}_{4}$
1 0 0 0.707 0 0 0 0 0
2 0 0 0.707 0 0 0 0 0
3 0 0 0 0 0 0 0 -0.577
4 0 0 0 0 0 0 0 -0.577
5 0 0 0 0 0 0 0 -0.577
6 0 -0.5 0 0 0 0 0 0
7 0 -0.5 0 0 0 0 0 0
8 0 -0.5 0 0 0 0 0 0
9 0 -0.5 0 0 0 0 0 0
10 -0.447 0 0 0 0 0 0 0
11 -0.447 0 0 0 0 0 0 0
12 -0.447 0 0 0 0 0 0 0
13 -0.447 0 0 0 0 0 0 0
14 -0.447 0 0 0 0 0 0 0
15 0 0 0 0 0 0 1 0
16 0 0 0 0 0 1 0 0
17 0 0 0 0 1 0 0 0
18 0 0 0 1 0 0 0 0
Table 1: Example of combinatorial Laplacian eigenvectors related to zero eigenvalue for a dataset consisting of 4 separated groups of documents (of sizes 2,3,4,5) with one document (left table) or four documents (right table) unrelated to anything
Let us mention also issues with normalized Laplacian $\mathcal{L}(S)=D(S)^{-1/2}L(S)D(S)^{-1/2}$. For the $S^{\prime}$ matrix described above, $\mathcal{L}(S^{\prime})$ cannot be computed as one of elements of $D(S^{\prime})$ is equal to zero, so the inverted square root cannot be computed. Therefore we need to investigate other scenarios. First, note that
$\mathcal{L}(S)=\begin{pmatrix}\mathcal{L}(S_{[[1]]})&\mathbf{0}&\mathbf{0}&\dots&\mathbf{0}\\ \mathbf{0}&\mathcal{L}(S_{[[2]]})&\mathbf{0}&\dots&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathcal{L}(S_{[[3]]})&\dots&\mathbf{0}\\ \dots&\dots&\dots&\dots&\dots\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\dots&\mathcal{L}(S_{[[k]]})\end{pmatrix}$, (27)
Consider now the following similarity matrix
$S^{**}=\begin{pmatrix}S&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&0&\epsilon\\ \mathbf{0}&\epsilon&0\end{pmatrix}=\begin{pmatrix}S_{[[1]]}&\mathbf{0}&\mathbf{0}&\dots&\mathbf{0}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&S_{[[2]]}&\mathbf{0}&\dots&\mathbf{0}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&S_{[[3]]}&\dots&\mathbf{0}&\mathbf{0}&\mathbf{0}\\ \dots&\dots&\dots&\dots&\dots&\dots&\dots\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\dots&S_{[[k]]}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\dots&\mathbf{0}&0&\epsilon\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\dots&\mathbf{0}&\epsilon&0\end{pmatrix}$, (28)
Then
$\mathcal{L}(S^{**)}=\mathcal{L}\begin{pmatrix}S&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&0&\epsilon\\ \mathbf{0}&\epsilon&0\end{pmatrix}=\begin{pmatrix}\mathcal{L}(S_{[[1]]})&\mathbf{0}&\mathbf{0}&\dots&\mathbf{0}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathcal{L}(S_{[[2]]})&\mathbf{0}&\dots&\mathbf{0}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&S_{[[3]]}&\dots&\mathbf{0}&\mathbf{0}&\mathbf{0}\\ \dots&\dots&\dots&\dots&\dots&\dots&\dots\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\dots&\mathcal{L}(S_{[[k]]})&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\dots&\mathbf{0}&\frac{\epsilon}{\sqrt{\epsilon\epsilon}}&-\frac{\epsilon}{\sqrt{\epsilon\epsilon}}\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\dots&\mathbf{0}&-\frac{\epsilon}{\sqrt{\epsilon\epsilon}}&\frac{\epsilon}{\sqrt{\epsilon\epsilon}}\end{pmatrix}$, (29)
$k+1$ eigenvalues of $\mathcal{L}(S^{**)}$ are equal zero whereby the first $k$ corresponding eigenvectors are those of $L(S)$ and the last one is of the form $(\mathbf{0},\frac{1}{sqrt{2}},\frac{1}{sqrt{2}})^{T}$. The argument of the problem with such unrelated documents goes here also along the already outlined argument for combinatorial Laplacian, but with higher cardinalities of clusters. The table 2 presents example eigenvectors obtained when adding 2 or 4 unrelated documents to a document collection with cleanly separated 4 clusters of 4,6,8,10 documents resp.
$\mathbf{o}$ $\mathbf{v}_{1}$ $\mathbf{v}_{5}$ $\mathbf{v}_{2}$ $\mathbf{v}_{3}$ $\mathbf{v}_{4}$
1 0 0 0 0.53 0
2 0 0 0 0.471 0
3 0 0 0 0.504 0
4 0 0 0 0.493 0
5 0 0 -0.411 0 0
6 0 0 -0.391 0 0
7 0 0 -0.41 0 0
8 0 0 -0.415 0 0
9 0 0 -0.422 0 0
10 0 0 -0.4 0 0
11 -0.352 0 0 0 0
12 -0.355 0 0 0 0
13 -0.359 0 0 0 0
14 -0.354 0 0 0 0
15 -0.351 0 0 0 0
16 -0.349 0 0 0 0
17 -0.358 0 0 0 0
18 -0.35 0 0 0 0
19 0 0 0 0 -0.322
20 0 0 0 0 -0.318
21 0 0 0 0 -0.314
22 0 0 0 0 -0.316
23 0 0 0 0 -0.308
24 0 0 0 0 -0.324
25 0 0 0 0 -0.315
26 0 0 0 0 -0.309
27 0 0 0 0 -0.322
28 0 0 0 0 -0.314
29 0 0.707 0 0 0
30 0 0.707 0 0 0
$\mathbf{o}$ $\mathbf{v}_{1}$ $\mathbf{v}_{5}$ $\mathbf{v}_{6}$ $\mathbf{v}_{2}$ $\mathbf{v}_{3}$ $\mathbf{v}_{4}$
1 0 0 0 -0.523 0 0
2 0 0 0 -0.501 0 0
3 0 0 0 -0.49 0 0
4 0 0 0 -0.486 0 0
5 0 0 0 0 -0.414 0
6 0 0 0 0 -0.419 0
7 0 0 0 0 -0.408 0
8 0 0 0 0 -0.387 0
9 0 0 0 0 -0.416 0
10 0 0 0 0 -0.406 0
11 -0.338 0 0 0 0 0
12 -0.345 0 0 0 0 0
13 -0.347 0 0 0 0 0
14 -0.362 0 0 0 0 0
15 -0.359 0 0 0 0 0
16 -0.358 0 0 0 0 0
17 -0.36 0 0 0 0 0
18 -0.359 0 0 0 0 0
19 0 0 0 0 0 -0.314
20 0 0 0 0 0 -0.317
21 0 0 0 0 0 -0.316
22 0 0 0 0 0 -0.324
23 0 0 0 0 0 -0.319
24 0 0 0 0 0 -0.323
25 0 0 0 0 0 -0.308
26 0 0 0 0 0 -0.313
27 0 0 0 0 0 -0.312
28 0 0 0 0 0 -0.314
29 0 0 0.707 0 0 0
30 0 0 0.707 0 0 0
31 0 0.707 0 0 0 0
32 0 0.707 0 0 0 0
Table 2: Example of normalized Laplacian eigenvectors related to zero eigenvalue for a dataset consisting of 4 separated groups of documents (of sizes 4,6,8,10) with one pair of documents (left table) or two pairs of documents (right table) unrelated to anything
To check the impact of noise, we generated a similarity matrix for 4 clusters with cluster sizes 8 12 16 20. Then added noise in the range from=0., to=0.000001. $k$-means (with 200fold restart) applied to the space spanned by 4 eigenvectors of the combinatorial Laplacian related to 4 lowest eigenvalues retrieved correctly 4 clusters with sizes 8 12 16 20.
However, for the same design, if one pair of unrelated documents was added, one got cluster sizes 20, 28, 8, 2.
For true clusters of sizes 16 24 32 40 the returned clusters, when 2 pairs of unrelated documents were added, were: 2, 66, 32, 16.
In all these examples we see that tiny one-doc or two-doc start to impact the final clustering badlyu. Their impct could be elevated if our method of removal of unwanted (unrelated) documents would be applied.
@@ 4.3 Real data
This fact causes the $L$-based and $K$-based and $\mathcal{B}$-based clustering algorithms to perform poorly, see Table 6, 5, 8. Compared to them, the $N$-based algorithm performs quite well, see Table 7.
@@
The tables 9, 10, 12, 11 show the impact of this filtering. In each case one can see an improvement of the performance.
@@ 4.5 Real Data Experiments - Multiple Clustering Methods
The Tables 15, …, 20 present results of our experiments about the impact on explainability when performing rough-set based removal of boundary data.